OpenStudy (anonymous):

Solve the equation 4 over 5 y + 2 = 6 for y. Please show your work and solution

OpenStudy (anonymous):

\[\frac{4}{5y+2}=6\] \[\frac{4}{6}=5y+2\] \[\frac{2}{3}=5y+2\] \[5y=\frac{2}{3}-2\] \[5y=\frac{2-6}{3}\] \[y=\frac{-4}{3*5}\] \[y=\frac{-4}{15}\]